PLEAS HELP
USE LINEAR COMBINATIONS TO SOLVE THE FOLLOWING
1. 2x -3y = -16
x + 3 = 10



2. 3x + y = -8
-3x + 4y = -2


3. 2x -y = 1
2x + 5y = -5

please i need some help.

Answer by venugopalramana(3286)

SEE THE FOLLOWING AND TRY.IF STILL IN DIFFICULTY PLEASE COME BACK
you gave me a site to go on and i can't get on it i am trying to see how to do these because i am out of school right now and hve no one to show me
use linear combinations to solve the system linera equation
THESE ARE ALL SIMILAR PROBLEMS.I SHALL SHOW ONE AND YOU MAY TRY THE OTHERS.
1. x + 2 =5....................EQN.I
3x - 2y =7.....................EQN.II
1.THERE ARE 2 UNKNOWNS AND 2 EQNS.
2.IF THERE IS 1 UNKNOWN AND 1 EQN.WE CAN SOLVE IT
3.SO WE TRY TO REMOVE ONE UNKNOWN BY ELIMINATION.
EQN.I GIVES X=5-2=3
EQN.II ..3*3-2Y=7
2Y=9-7=2
Y=1
2. x + y = 1......................III
2x -3y = 12........................IV
EQN.I*2-EQN.II
2X+2Y-2X+3Y=2-12=-10
5Y=-10
Y=-2
PUT THIS IN EQN.III
X-2=1
X=3....ETC..
3. x -y = -4
x + 2y = 5